# Markov Perfect Equilibria in

### v1

@@ 2.3 The mean-field problem
A mean-field game (Nash) equilibrium (MFG-NE), started from $\mu\in\Sigma^{d-1}$ at time $t\in{\mathcal{T}}$, is a measure flow $\boldsymbol{\mu}^{*}=(\mu^{*}_{t},\ldots,\mu^{*}_{T})$ such that there exists a control $\alpha^{*}\in\mathscr{A}$ satisfying the following.
@@
    Consistency: For all $t\leq s\leq T$, we have
    $\mu^{*}_{s}(x)=\sum_{y\in{\mathcal{X}}}\mu(y)\,\mathbb{P}^{\alpha^{*}}[X_{s}=x\,|\,X_{t}=y],\quad x\in{\mathcal{X}}.$
@@
It is well known, see Remark 2.15 below, that Nash equilibria correspond to solutions of the forward-backward mean-field game system consisting of dynamic programming and Kolmogorov equations for $x\in{\mathcal{X}}$,
$\left\{\begin{split}v(s,x)&=\underset{a\in A(h,x)}{\inf}\,H(x,\mu_{s},v(s+h),a),\quad t\leq s\leq T,\\ \alpha(s,x)&=\underset{a\in A(h,x)}{\mathrm{arg\,min}}\ H(x,\mu_{s},v(s+h),a),\quad t\leq s<T,\\ \mu_{s+h}(x)&=\sum_{y\in{\mathcal{X}}}\mu_{s}(y)\alpha_{x}(s,y),\quad t\leq s<T,\\ \mu_{t}&=\mu,\ v(T,x)=g(x,\mu_{T}).\end{split}\right.$ (MFG system)
@@
More precisely, under Assumption 2.5 and given initial data $(t,\mu)\in{\mathcal{T}}\times\Sigma^{d-1}$, a measure flow $\boldsymbol{\mu}\in(\Sigma^{d-1})^{{\mathcal{T}}(t,T)}$ is a MFG-NE starting from initial distribution $\mu$ at time $t$ if and only if there exists a function $v:{\mathcal{T}}(t,T)\times{\mathcal{X}}\to[0,\infty)$ such that $(v,\boldsymbol{\mu})$ solve the MFG system with initial data $(t,\mu)$.
@@
[…]
@@
Let Assumptions 2.5 hold and let $\bar{\alpha}\in\mathscr{A}_{\infty}$. Then $\bar{\alpha}$ is a MF-MPE if and only if there exists a function $V:\bar{\mathcal{T}}\times\mathscr{X}\to[0,\infty)$ such that $(V,\bar{\alpha})$ satisfies (MF-NLL).
@@
    There exist a solution to the mean-field game starting from $\mu$ at time $t$.
@@
    Let $(V,\bar{\alpha})$ be a solution to (MF-NLL). Then, the flow $(\mu^{*}_{s})_{t\leq s\leq T}$, defined by
    $\mu^{*}_{t}:=\mu,\quad\mu^{*}_{s+h}(x):=\sum_{y\in{\mathcal{X}}}\mu^{*}_{s}(y)\bar{\alpha}_{x}(s,y,\mu^{*}_{s}),\quad v^{*}(s,x):=V(s,x,\mu^{*}_{s}),$
    is a solution to the MFG system.
The implications in Proposition 2.18 can be illustrated as follows.
[…]
@@ 2.4 Short-time results
Let Assumptions 2.5, 2.7 and 2.19 hold. Then there exist constants $h_{*},T_{*}>0$, depending on the data and $m=m(d)$ but not on the number of players $N+1$, such that, whenever $h<h_{*}$ and $Kh<T_{*}$, there exists a unique solution to (MF-NLL) and ($N$-NLL) on $\bar{\mathcal{T}}=\{0,h,\ldots,Kh\}$, for any $N\geq 1$.
@@ 2.5 Mean-field convergence
Let the assumptions of Theorem 2.20 hold, and let $h_{*},T_{*}$ be as in that theorem. Let $h<h_{*}$ and $Kh<T_{*}$ be given. Let $(v^{(N)},\alpha^{(N)})$ solve ($N$-NLL) and let $(V,\alpha)$ be the solution to (MF-NLL). Set $\bar{\mathcal{T}}=\{0,h,\ldots,Kh\}$. Then the following hold:
@@ 2.6 Continuous-time limit
We now fix the time horizon $T$ and the number of players $N+1$, and we consider the continuous-time limit $h\downarrow 0$. We continue to briefly define the limiting problem and refer to [31] for a detailed discussion. In the continuous-time model, players’ state processes follow a time-inhomogeneous continuous-time Markov chain, which, by an abuse of notation, we again denote by $(X_{t},Z_{t})$. For a given thermal noise level $\sigma^{2}\geq 0$, let $\beta\in\mathscr{A}_{\mathrm{cts}}$ be an admissible feedback control, that is, a Lipschitz continuous function $\beta:[0,T]\times\mathscr{X}_{N}\to\mathbb{R}^{d}$ satisfying, for any $(t,x,z)\in[0,T]\times\mathscr{X}_{N}$,
$\beta_{x^{\prime}}(t,x,z)\geq\sigma^{2}\quad\text{for all}\ x^{\prime}\in{\mathcal{X}}\setminus\{x\}\ \text{and}\quad\beta_{x}(t,x,z)=-\sum_{y\neq x}\beta_{y}(t,x,z).$
Here, $\beta_{x^{\prime}}(t,x,z)$ is the rate of changing from the current state $x\in{\mathcal{X}}$ to a state $x^{\prime}\neq x$ at time $t\in[0,T]$, given that the empirical distribution of other players is $z\in\Sigma^{d-1}_{N}$. Given a control $\alpha\in\mathscr{A}_{\mathrm{cts}}$ of the tagged player and a control $\beta\in\mathscr{A}_{\mathrm{cts}}$ used by all other players, for any $(t,x,z)\in[0,T]\times\mathscr{X}_{N}$ and $y,w,x^{\prime}\in{\mathcal{X}}$, we have
@@
where we recall the definition $e^{N}_{xy}=\frac{1}{N}(e_{x}-e_{y})$ and $o(\Delta t)/\Delta t\to 0$ as $\Delta t\downarrow 0$. The tagged player solves
@@
To define the NLL equation, we make the following explicit assumption.
Assumption 2.22.
For every $(x,\mu,a)\in\mathscr{X}\times[0,\infty)^{d}$, the running cost $\ell(x,\mu,a)$ is independent of $a_{x}$.
Given $\beta:\mathscr{X}_{N}\to\mathbb{R}^{d}$, we define the generator $\mathscr{L}^{\beta}$ of the process $Z_{t}$ by
$(\mathscr{L}^{\beta}\varphi)(x,z):=\sum_{w\neq y,\,w,y\in{\mathcal{X}}}Nz_{y}\beta_{w}(y,z+e^{N}_{xy})[\varphi(x,z+e^{N}_{wy})-\varphi(x,z)],\quad(x,z)\in\mathscr{X}_{N},$
@@
${\mathcal{H}}(x,z,v,a):=\ell(x,z,a)+a\cdot v,\quad(x,z,v,a)\in\mathscr{X}_{N}\times\mathbb{R}^{d-1}\times[0,\infty)^{d-1}.$
Finally, we denote the minimizer of the Hamiltonian by $\hat{\alpha}:\mathscr{X}_{N}\times\mathbb{R}^{d-1}\to\mathbb{R}^{d}$, i.e.,
$(\hat{\alpha}_{y}(x,z,v))_{y\in{\mathcal{X}}\setminus\{x\}}:=\underset{a\in[\sigma^{2},\infty)^{d-1}}{\mathrm{arg\,min}}\ {\mathcal{H}}(x,z,v,a)\quad\text{and}\quad\hat{\alpha}_{x}(x,z,v):=-\sum_{y\neq x}\hat{\alpha}_{y}(x,z,v).$
Gomes, Mohr and Souza [31] show that there is a unique MPE which corresponds to the unique classical solution $v_{\mathrm{cts}}:[0,T]\times\mathscr{X}_{N}\to[0,\infty)$ of the following continuous-time finite-player NLL equation,
$\left\{\begin{split}-\frac{\mathrm{d}}{\mathrm{d}t}v_{\mathrm{cts}}(t,x,z)&=\underset{a\in[\sigma^{2},\infty)^{d-1}}{\inf}{\mathcal{H}}(x,z,\Delta_{x}v_{\mathrm{cts}}(t,\cdot,z),a)+\mathscr{L}^{\alpha_{\mathrm{cts}}(t)}v(t)(x,z),\\ \alpha_{\mathrm{cts}}(t,x,z)&=\hat{\alpha}(x,z,\Delta_{x}v_{\mathrm{cts}}(t,\cdot,z)),\\ v_{\mathrm{cts}}(T,x,z)&=g(x,z).\end{split}\right.$ (2.9)
@@
$\Delta_{x}\varphi:=(\varphi(y)-\varphi(x))_{y\neq x}\in\mathbb{R}^{d-1}$
for any $\varphi:{\mathcal{X}}\to\mathbb{R}$. We continue to state our convergence result. For $K\geq 1$, define $h=h^{(K)}:=T/K$ and $\bar{\mathcal{T}}^{(K)}:=\{0,h,\ldots,(K-1)h,T\}$ and $A(h)=A(h,x)=\{a\in\Sigma^{d-1}:a_{y}\geq h\sigma^{2}\ \forall y\in{\mathcal{X}}\}$. For sufficiently large $K$, let $(v^{(K)},\alpha^{(K)})$ be the unique solution to the ($N$-NLL) equation on $\bar{\mathcal{T}}^{(K)}$, with the aforementioned choice of $A(h)$, and extend $(v^{(K)},\alpha^{(K)})$ to $[0,T]\times\mathscr{X}_{N}$ by linear interpolation in the time variable. Set
@@
Under Assumptions 2.5, 2.7, 2.19 and 2.22, we have the following.
@@
    (Convergence of values) As $K\to\infty$, $v^{(K)}\to v_{\mathrm{cts}}$ uniformly on $[0,T]\times\mathscr{X}_{N}$.
@@
    (Convergence of MPE) As $K\to\infty$, $\mathfrak{a}^{(K)}/h^{(K)}\to\alpha_{\mathrm{cts}}$ uniformly on $[0,T]\times\mathscr{X}_{N}$.
@@ 2.7 Comparison with existing works
In Remarks 2.6 and 2.17, we first clarify the connection between symmetric Markov perfect equilibria and the Nash-Lasry-Lions equation, both in the finite-player and mean-field formulations. In the finite-player case, this relationship has been observed, for example by Gomes, Mohr and Souza [31] for a reference close to our setting.
Next, we address the existence and uniqueness of MPE in our finite-player and mean-field problems, see Theorem 2.10. Existence follows from well-known fixed-point arguments, see for example Gomes, Mohr and Souza [30], while our uniqueness proof relies on a contraction argument and uses the smallness of the time step. This result sheds light on the relationship with the continuous-time problem, which in the finite-player case becomes uniquely solvable, see [31]. In the mean-field, we obtain short-time results in Theorem 2.20, which can be made global under monotonicity assumptions. In this regard, our work is related to the pasting of equilibria by Dianetti, Nendel, Tangpi and Wang [23], who utilize a discrete-time master equation approach under Lasry-Lions monotonicity, see Remark 4.2 below.
Finally, we present convergence results of the finite-player game in Theorems 2.21 and 2.23, both as the number of players $N+1$ goes to infinity and as the time discretization $h$ goes to zero. The former parallels the results obtained in [31] in continuous-time while in the latter limit we recover the uniquely solvable finite-player NLL equation in [31].
Structure. The rest of the paper is structured as follows. In Section 3 we investigate the symmetric finite-player problem and establish Theorem 2.10. Section 4 discusses the mean-field problem and proves Proposition 2.18. The short-time results recorded in Theorem 2.20 for both finite-player and mean-field games are obtained in Section 5. The mean-field limit $N\to\infty$, see Theorem 2.21, is presented in Section 6, and the continuous-time limit $h\downarrow 0$, see Theorem 2.23, is treated in Section 7. We recall some basic facts about dynamic programming in discrete time in Appendix A.1 and prove a generalized conjugate correspondence theorem in Appendix A.2.
@@ 3 The symmetric finite-player problem
This section proves Theorem 2.10. While the proof of (i) rely on results in dynamic programming and an application of Brouwer’s fixed-point theorem, the analysis for part (ii) makes crucial use of a Lipschitz property of the multinomial distribution presented in Lemmas 3.1, 3.2.

### v2

@@ 2.3 The mean-field problem
A solution to the mean-field game or mean-field game (Nash) equilibrium (MFG-NE), started from $\mu\in\Sigma^{d-1}$ at time $t\in{\mathcal{T}}$, is a measure flow $\boldsymbol{\mu}^{*}=(\mu^{*}_{t},\ldots,\mu^{*}_{T})$ such that there exists a control $\alpha^{*}\in\mathscr{A}$ satisfying the following.
@@
    Consistency: For all $t\leq s\leq T$, we have μ^*_s(x) = ∑_y∈X μ(y)   P^α^*[X_s=x | X_t=y], x∈X.
@@
It is well known, see Remark 2.16 below, that Nash equilibria correspond to solutions of the forward-backward mean-field game system consisting of dynamic programming and Kolmogorov equations for $x\in{\mathcal{X}}$ and $t\leq s<T$:
$\left\{\begin{split}v(s,x)&=\underset{a\in A(h,x)}{\inf}\,H(x,\mu_{s},v(s+h),a),\\ \alpha(s,x)&=\underset{a\in A(h,x)}{\mathrm{arg\,min}}\ H(x,\mu_{s},v(s+h),a),\\ \mu_{s+h}(x)&=\sum_{y\in{\mathcal{X}}}\mu_{s}(y)\alpha_{x}(s,y),\\ \mu_{t}&=\mu,\ v(T,x)=g(x,\mu_{T}).\end{split}\right$. (MFG system)
@@
More precisely, under Assumption 2.6 and given initial data $(t,\mu)\in{\mathcal{T}}\times\Sigma^{d-1}$, a measure flow $\boldsymbol{\mu}\in(\Sigma^{d-1})^{{\mathcal{T}}(t,T)}$ is a MFG-NE starting from initial distribution $\mu$ at time $t$ if and only if there exists a function $v:{\mathcal{T}}(t,T)\times{\mathcal{X}}\to[0,\infty)$ such that $(v,\boldsymbol{\mu})$ solve (MFG system) with initial data $(t,\mu)$.
@@
$\left\{\begin{aligned} V(t,x,\mu)&=\underset{a\in A(h,x)}{\inf}\,H(x,\mu,V(t+h,\cdot,\mu\cdot\bar{\alpha}(t,\cdot,\mu)),a),\\ \bar{\alpha}(t,x,\mu)&=\underset{a\in A(h,x)}{\mathrm{arg\,min}}\,H(x,\mu,V(t+h,\cdot,\mu\cdot\bar{\alpha}(t,\cdot,\mu)),a),\\ V(T,x,\mu)&=g(x,\mu).\end{aligned}\right$. (MF-NLL)
@@
Let Assumptions 2.6 hold and let $\bar{\alpha}\in\mathscr{A}_{\infty}$. Then, $\bar{\alpha}$ is a MF-MPE if and only if there exists a function $V:\bar{\mathcal{T}}\times\mathscr{X}\to[0,\infty)$ such that $(V,\bar{\alpha})$ satisfies (MF-NLL).
@@
    There exists a solution to the mean-field game starting from $\mu$ at time $t$.
@@
    Let $(V,\bar{\alpha})$ be a solution to (MF-NLL). Then, the pair $(v^{*},\mu^{*})$, defined by μ^*_t := μ, μ^*_s+h(x) := _yX μ^*_s(y)¯α_x(s,y,μ^*_s), v^*(s,x):=V(s,x,μ^*_s),  st<T, is a solution to (MFG system).
The implications above can be illustrated as follows.
MF-MPEsolutions to MF-NLLsolutions to MFG systemMFG-NE
@@ 2.4 Short-time results
Let Assumptions 2.6, 2.8 and 2.20 hold. Then, there exist constants $h_{*},T_{*}>0$, depending on the data and $m$ but not on the number of players $N+1$, such that, whenever $h<h_{*}$ and $Kh<T_{*}$, there exists a unique solution to (MF-NLL) and (NN-NLL) on $\bar{\mathcal{T}}=\{0,h,\ldots,Kh\}$, for any $N\geq 1$.
@@ 2.5 Mean-field convergence
Let the assumptions of Theorem 2.21 hold, and let $h_{*},T_{*}$ be as in that theorem. Let $h<h_{*}$ and $Kh<T_{*}$ be given. Let $(V,\alpha)$ and $(v^{(N)},\alpha^{(N)})$ be the solutions to (MF-NLL) and (NN-NLL), respectively. Set $\bar{\mathcal{T}}=\{0,h,\ldots,Kh\}$. Then the following hold:
@@ 2.6 Continuous-time limit
We now fix the time horizon $T$ and the number of players $N+1$, and we consider the continuous-time limit $h\downarrow 0$. We continue to briefly define the limiting problem and refer to [31] for a detailed discussion. In the continuous-time model, players’ state processes follow a time-inhomogeneous continuous-time Markov chain, which, by an abuse of notation, we again denote by $(X_{t},Z_{t})$. For a given thermal noise level $\sigma^{2}\geq 0$, let $\mathscr{A}_{\mathrm{cts}}$ be the set of all admissible feedback controls $\beta$, that is, Lipschitz continuous functions $\beta:[0,T]\times\mathscr{X}_{N}\to\mathbb{R}^{d}$ satisfying $\beta(t,x,z)\in{\mathcal{R}}(\sigma^{2},x)$ for any $(t,x,z)\in[0,T]\times\mathscr{X}_{N}$. Here, the set of admissible transition rates from a state $x$ is defined by
${\mathcal{R}}(\sigma^{2},x):=\Big\{a\in\mathbb{R}^{d}:a_{y}\geq\sigma^{2}\ \text{for all}\ y\in S(x)\setminus\{x\},\ a_{y}=0\ \text{for all}\ y\in{\mathcal{X}}\setminus(S(x)\cup\{x\})\ \text{and}\ a_{x}=-\sum_{y\neq x}a_{y}\Big\},$
where, as before, $S(x)\subset{\mathcal{X}}$ denotes a (fixed) set of allowable transitions. In words, $\beta_{y}(t,x,z)$ is the rate of changing from the current state $x\in{\mathcal{X}}$ to a state $y\neq x$ at time $t\in[0,T]$, given that the empirical distribution of other players is $z\in\Sigma^{d-1}_{N}$. Given a control $\alpha\in\mathscr{A}_{\mathrm{cts}}$ of the tagged player and a control $\beta\in\mathscr{A}_{\mathrm{cts}}$ used by all other players, for any $(t,x,z)\in[0,T]\times\mathscr{X}_{N}$ and $y,w,x^{\prime}\in{\mathcal{X}}$, we have
@@
where we recall the definition $e^{N}_{xy}=\frac{1}{N}(e_{x}-e_{y})$, and $o(\Delta t)/\Delta t\to 0$ as $\Delta t\downarrow 0$. The tagged player solves
@@
Given $\beta:\mathscr{X}_{N}\to\mathbb{R}^{d}$, we define the operator $\mathscr{L}^{\beta}$ by setting
$(\mathscr{L}^{\beta}\varphi)(x,z):=\sum_{w,y\in{\mathcal{X}}:\,w\neq y}Nz_{y}\beta_{w}(y,z+e^{N}_{xy})[\varphi(x,z+e^{N}_{wy})-\varphi(x,z)],\quad(x,z)\in\mathscr{X}_{N},$ (2.9)
@@
${\mathcal{H}}(x,z,v,a):=\ell(x,z,a)+a\cdot v,\quad(x,z,v,a)\in\mathscr{X}_{N}\times\mathbb{R}^{d}\times\mathbb{R}^{d}.$
Finally, we denote the minimizer of the Hamiltonian by
$\hat{\alpha}:\mathscr{X}_{N}\times\mathbb{R}^{d}\to\mathbb{R}^{d},\qquad\hat{\alpha}(x,z,v):=\underset{a\in{\mathcal{R}}(\sigma^{2},x)}{\mathrm{arg\,min}}\ {\mathcal{H}}(x,z,v,a).$
Gomes, Mohr and Souza [31] show that there is a unique MPE which corresponds to the unique classical solution $v_{\mathrm{cts}}:[0,T]\times\mathscr{X}_{N}\to[0,\infty)$ of the following continuous-time finite-player NLL equation:
$\left\{\begin{split}-\frac{\mathrm{d}}{\mathrm{d}t}v_{\mathrm{cts}}(t,x,z)&=\underset{a\in{\mathcal{R}}(\sigma^{2},x)}{\inf}{\mathcal{H}}(x,z,\Delta_{x}v_{\mathrm{cts}}(t,\cdot,z),a)+\mathscr{L}^{\alpha_{\mathrm{cts}}(t)}v(t)(x,z),\\ \alpha_{\mathrm{cts}}(t,x,z)&=\hat{\alpha}(x,z,\Delta_{x}v_{\mathrm{cts}}(t,\cdot,z)),\\ v_{\mathrm{cts}}(T,x,z)&=g(x,z).\end{split}\right$. (2.10)
@@
$\Delta_{x}\varphi:=(\varphi(y)-\varphi(x))_{y\in{\mathcal{X}}}\in\mathbb{R}^{d}$
for any $\varphi:{\mathcal{X}}\to\mathbb{R}$. Note that $\Delta_{x}\varphi(x)=0$. We continue to state our convergence result. For $K\geq 1$, define $h=h^{(K)}:=T/K$ and $\bar{\mathcal{T}}^{(K)}:=\{0,h,\ldots,(K-1)h,T\}$ and we assume
$A(h,x)=\big\{a\in\Sigma^{d-1}:a_{y}\geq h\sigma^{2}\ \text{for all}\ y\in S(x)\setminus\{x\}\ \text{and}\ a_{y}=0\ \text{for all}\ y\in{\mathcal{X}}\setminus(S(x)\cup\{x\})\big\}.$ (2.11)
For sufficiently large $K$, let $(v^{(K)},\alpha^{(K)})$ be the unique solution to ($N$-NLL) on $\bar{\mathcal{T}}^{(K)}$, guaranteed by Theorem 2.11, with the aforementioned choice of $A(h,x)$, and extend $(v^{(K)},\alpha^{(K)})$ to $0,T]\times\mathscr{X}_{N}$ by linear interpolation in the time variable. Set
@@
Under Assumptions 2.6, 2.8 and 2.20 we have the following.
@@
    (Convergence of values) As $K\to\infty$, $v^{(K)}\to v_{\mathrm{cts}}$ uniformly on $0,T]\times\mathscr{X}_{N}$.
@@
    (Convergence of MPE) As $K\to\infty$, $\mathfrak{a}^{(K)}/h^{(K)}\to\alpha_{\mathrm{cts}}$ uniformly on $0,T]\times\mathscr{X}_{N}$.
@@ 2.7 Comparison with existing works
In Remarks 2.7 and 2.18, we first clarify the connection between symmetric Markov perfect equilibria and the Nash-Lasry-Lions equation, both in the finite-player and the mean-field formulation. In the finite-player case, this relationship has been observed, see for example Gomes, Mohr and Souza [31] for a reference close to our setting.
Next, we address the existence and uniqueness of MPE in our finite-player problem, see Theorem 2.11. Existence follows from well-known fixed-point arguments, see for example Gomes, Mohr and Souza [30], while our uniqueness proof relies on a contraction argument and uses the smallness of the time step. This result sheds light on the relationship with the continuous-time problem, which in the finite-player case becomes uniquely solvable, see [31]. In the mean-field, we obtain short-time results in Theorem 2.21, which can be made global under monotonicity assumptions. In this regard, our work is related to the pasting of equilibria by Dianetti, Nendel, Tangpi and Wang [23], who utilize a discrete-time master equation approach under Lasry-Lions monotonicity, see Remark 4.2 below.
Finally, we present convergence results of the finite-player game in Theorems 2.22 and 2.23, both as the number of players $N+1$ goes to infinity and as the time discretization $h$ goes to zero. The former parallels the mean-field convergence obtained in [31] in continuous-time while in the latter limit we recover the uniquely solvable finite-player NLL equation in [31].
Structure. The rest of the paper is organized as follows. In Section 3 we investigate the symmetric finite-player problem and establish Theorem 2.11. Section 4 discusses the mean-field problem and proves Proposition 2.19. The short-time results recorded in Theorem 2.21 for both finite-player and mean-field games are obtained in Section 5. The mean-field limit $N\to\infty$, see Theorem 2.22, is established in Section 6, and the continuous-time limit $h\downarrow 0$, see Theorem 2.23, is treated in Section 7. We recall some basic facts about dynamic programming in discrete time in Appendix A.1 and prove a generalized conjugate correspondence theorem in Appendix A.2.
@@ 3 The symmetric finite-player problem
This section proves Theorem 2.11. While the proof of (i) relies on results in dynamic programming and an application of Brouwer’s fixed-point theorem, the analysis for part (ii) makes crucial use of a Lipschitz property of the multinomial distribution presented in Lemmas 3.1, 3.2.